Planks Plants had net income of $5000 on sales of $90.000 last year The firm paid a dividend of $1.550. Total assets were $300.000 of which $150,000 was financed by debt a. What is the firms sustainable growth rate? (Do not round intermediate calculations. Enter your answer as a percent rounded to 1 decimal place.) Sustainable growth rate b. If the firm grows at its sustainable growth rate, how much debt will be issued next year? (Do not round intermediate calculations.) New debt c. What wouild be the meximum possibie growth rate if the firm did not issue any debt next year? (Do not round intermediate calculations. Enter your answer as e percent rounded to 1 decimal place.) Maximum growth rate

I attempted and the system said I was wrong. where am I messing up?.
Answer #1

1) Sustainable growth rate = 6.9%

Working: (5000 - 1550) / 5,000 * 5000 / 50,000 = 6.9%

2) New debt = 10,350

Working: 300,000 * 6.9% = 20,700

150,000 / 300,000 = 0.5

0.5 * 20,700 = 10,350

3) Maximum growth rate = 3.45%

Working: (5000 - 1550) / 5,000 * 5000 / 50,000 * 0.5% = 3.45%